OpenStudy (anonymous):

Find the range of the function rule y = 3x + 4 for each domain.

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

9. {2, 9, 11.5}

OpenStudy (anonymous):

@nirmalnema

OpenStudy (nirmalnema):

all real nos.. @Ozymandios is correct

OpenStudy (anonymous):

no it is not all real numbers

OpenStudy (anonymous):

the rule is \[y = 3x + 4\] and if the domain is \[\{2, 9, 11.5 \}\] then the range is what you get when you plug these three numbers in to the formula

OpenStudy (anonymous):

\[3\times 2+4=10,3\times 9+4=31, 3\times 11.5+4=38.5\] so your range is \[\{10,31,38.5\}\]

OpenStudy (anonymous):

Ty ! Whats (-3,0,1,4)?

OpenStudy (anonymous):

all real numbers was a reflex action from not reading the entire question you have to plug the numbers in and see what you get

OpenStudy (anonymous):

i am sure you can compute the four numbers that would be the range plug in \(-3,0,1,\) and \(4\) in to the formula \(y=3x+4\) and see what you get i.e take each number, multiply it by 3, add 4 and see what you get

OpenStudy (anonymous):

I got { -5 , 4 ,7 , 16} is that right?

OpenStudy (anonymous):

@satellite73